Manuel Kauers Hagenberg, 19. März 2003 RISC-Linz

Word Order in German Sentences

Main Clauses and Subclauses II

Recall

Simple Statements, Questions, and Imperatives

• Sentences consist of building blocks which can be arranged in several ways to form a sentence.

• Every sentence has at least the subject and the verb. Everything but the verb may be put in an arbitrary order, but don’t put the subject to far to the end.

• The verb usually goes to the second position. If it splits into parts, then the conjugated part goes to the second position and the non conjugated part goes to the end:

teilnehmen – participate

Er nimmt am Deutschkurs teil.

• Simple questions (questions without questioning word) have the conjugated part of the verb at the first position.

Nimmt er am Deutschkurs teil?

• Imperatives are formed like questions:

nächste – next Woche, f – week

Nimm nächste Woche am Deutschkurs teil!

Composing clauses to sentences

• A simple clause (see above) may count as a sentence. Also, sentences can be made up of several clauses using connectives.

• A connective determines if it starts a main clause (Hauptsatz, m) or a subclause (Nebensatz, m).

• The main clause connectives are the following:

und and  oder or
aber but  denn since
Nicht (nur) …, sondern (auch) … Not (only) …, but (also) …
Weder …, noch …  Neither …, nor …

Entweder …, oder … • Weder, Entweder, Nicht (nur) in the beginning may or may not count as

building blocks of the first sentence.

• The subclause connectives are the following:

weil/da because  obwohl  although
wenn  if (cond.)  ob  if, whether
während while  nachdem after
bevor  before  dass/daß that
als  as (temp.)  falls  in case
sofern  as long as  damit  for that
so daß  such that  indem  by

• In subclauses, the verb goes to the end. Composed verbs don’t split.

• Subclauses count as building block of the surrounding clause. They might be put at the first or last position, or after the verb.

Questions

• Simple questions are composed by putting the verb to the beginning of

the sentence:

sprechen – speak

Spricht Frau Mayr englisch?

• If a question word is used, this word comes the the first position and the verb goes to the sencond.

• Here are some question words:

wer  who  wo  where
wann  when  wieso/weshalb/warum why
wie  how  was  what
worüber on what  woher  from where

…

The corresponding default answers are

der/dieser this (one)  da  there
dann  then  deshalb/darum for this reason
so  so  das/dieses  this
darüber  on that  daher  from there

…

• Examples:

Bahnhof, m – railway station halten – (here) give über – (here) on daran – at this schwer – difficult Angst, f – angst

Wo ist der Bahnhof? Wann beginnt der Deutschkurs? Wer hält die Vorlesung über Computer Algebra? Wie geht das? Was ist daran so schwer? Wovor hast Du Angst?

• Question marks can be more complicated. They may involve prepositions

like in the following examples.

viele – many bekommen – get am Tag – per day Seit – since wohnen – live/stay

Wie viele Mails bekommst Du am Tag? Von wem bekommst Du die meisten Mails? Seit wann wohnst Du in Österreich?

Exercise

• Besides the question words above, there are further question words which are build by wo (+ r) + preposition, like wovor and worüber in the table (the r is put to avoid vocal clashes). Guess the question word for “for what” (that’s too easy). Try to complete the question word from the prepositions that you know.

• The question word wer has to be declinated, depending on what you are asking for. The declination is analogous to the article der (Singular only). Write it down in a table and form a question sentence for each case.

• Translate to german:

far – weit

away – here: entfernt laptop – Laptop, m take place – stattfinden talk – Vortrag, m about – über

How far is Linz (away)? Do you have a laptop? When does the seminar take place? What is your talk about?

• Translate to english:

können – can man – ‘one’ Brot, n – bread kaufen – buy öffnen – open fahren – go (by car) jemand – anybody heute – today zum = zu + dem

Wo kann man in Hagenberg Brot kaufen? Wann öffnet der Unimarkt? Fährt jemand heute zum Billa?

Indirect Questions Clauses

• The question word from the previous section may also serve as connectives which introduce subclauses.

• These subclauses are called indirect question clauses.

• They correspond to a special case of the english relative clauses, for instance “I don’t know who will come.”

• Here is a table with main clauses that are commonly used in connection

with indirect question clauses:

Ahnung, f – anticipation gerne – willingly egal – equal denken – think

Ich habe keine Ahnung, … I have no idea
Ich wüßte gerne, …  I would (really) like to know
Können Sie mir sagen, … Do you know (formally)
Weißt Du, …  Do you know (informally)
Es ist egal, …  It doesn’t matter
Es ist mir egal, …  I don’t care
Ich kann mir denken, …  I can imagine

… • Examples:

Bahnhof, m – station schwierig – difficult fast – almost stehen – stand anrufen – call

Können Sie mir sagen, wo der Bahnhof ist? Ich wüßte gerne, warum Deutsch so schwierig ist. Es ist fast egal, wo im Satz das Subjekt steht. Es ist nicht egal, wo im Satz das Verb steht. Ich sage Dir nicht, wen ich anrufe.

• Question words that are made up of wo and a preposition split into parts

if they have to be in a case different from nominative.

Blume, f – flower pflücken – pick

Weißt Du, für wen er die Blumen pflückt?

• Indirect question sentences may also serve as subjects/objects, or other

adverbial parts of the surrounding sentence.

zu spät – late

bestrafen – punish Leben, n – life betten – to lie down liegen – to lie andern – others Grube – hole graben – dig fallen – fall hinein – into (it)

Wer zu spät kommt, den bestraft das Leben.^{1} Wie man sich bettet, so liegt man.^{2} Wer anderen eine Grube gräbt, fällt selbst hinein.^{3}

Exercise

• Convert the following questions to indirect ones using a phrase from the

table as main clause.

was – what machen – do wohnen – live/stay gehen – go gleich – soon zum – to the Seminar, n – seminar

Was machst Du morgen? Wo wohnst Du? Wer geht gleich zum Unimarkt? Ist morgen Seminar?^{4}

• The German philosopher Georg Schimmel once said

gebildet – educated wissen – know finden – find

Gebildet ist, wer weiss, wo er findet, was er nicht weiss.

Do you understand?

Relative Clauses

• Relative clauses are similar to indirect question clauses. Here the articles and the question word welcher (which) serve as subjunctions.

• They are declinated to the case which they have to have wrt. the sub clause verb. The declination differs to the declination of articles only in genitive singular dessen/deren/dessen instead of des/der/des and dative plural: denen instead of den.

• Their genter and number is according to the word which they refer to.

1Gorbatschow 2German saying 3German saying 4Hint: use ob

• Relative clauses are subclauses; the verb is put at the end.

• Relative clauses are put directly behind the word to which they refer.

• If there is no ambiguity, they may also be put at the end of the clause, if this makes the clause more readable.

• Examples:

Hund, m – dog bellen – bark beißen – bite Problem, n – problem seit – since/for Woche, f – week studieren – study völlig – completly trivial – trivial Aufgabe, f – exercise geben – give machen – make/do müssen – must zu – too einfach – easy

Hunde, die bellen, beißen nicht.^{5} Das Problem, das Du seit Wochen studierst, ist völlig trivial. Die Aufgaben, die wir im Deutschkurs, den Manuel gibt, machen müssen, sind zu einfach.

Exercise

• For each case build an example sentence with the relative pronoun in that case.

• Fill in the correct relative pronouns in the slots.

Folge, f – sequence beschränkt – bounded monoton – monotonous konvergieren – converge Witz, m – joke erzählen – tell langweilig – boring Kind, n – child spielen – play heißen – being called Satz, m – sentence auf – on Seite – page stehen – stand verstehen – understand

Folgen, . . . . . . beschränkt und monoton sind, konvergieren. Der Witz, . . . . . . Du erzählst, ist langweilig. Das Kind, . . . . . . . . . Du spielst, heißt Anna. (Caution!) Den Satz, . . . . . . auf Seite 1 steht, verstehe ich nicht.

• In English there is no distinction between implicit question clauses and relative clauses, both start with who, which, whose, etc. What is the difference in German? Give examples for both kind of clauses.

Special Constructions

• There are some special constructions that don’t fit to any of the constructions above. They are listed here.

• A table with approximate translations

um … zu X-en for X-ing …
je … desto …  the… the …
ohne … zu X-en without X-ing …

…

• Examples:

lernen – learn dort – there promovieren – doing PhD lernen – learn lügen – ly Wimper, f – eyelash

Er kommt nach Österreich, um Deutsch zu lernen. Viele kommen zum RISC, um dort zu promovieren. Je mehr Du lernst, desto mehr kannst Du. Er kann lügen, ohne mit der Wimper zu zucken.

5German saying.